40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 24 03:58:13 2012, kiwamu, Update, LSC, Y arm + PRMI part II 
    Reply  Fri Feb 24 15:01:31 2012, kiwamu, Update, LSC, Y arm + PRMI part II time_series.png
       Reply  Fri Feb 24 19:18:28 2012, kiwamu, Update, LSC, Y arm + PRMI : how they should look like expected_time_series.png
    Reply  Fri Feb 24 18:37:13 2012, rana, Update, LSC, Y arm + PRMI part II 
Message ID: 6310     Entry time: Fri Feb 24 03:58:13 2012     Reply to this: 6313   6315
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: Y arm + PRMI part II 

I tried the Yarm + PRMI configuration again.

The PRMI part was locked, but it didn't stay locked during the Y arm was brought to the resonance point.

I will post the time series data later.

 

(locking of the PRMI part)

Tonight I was able lock the PRMI when the arm was off from the resonance by 10 nm (#6306).

This time I used REFL11Q to lock the MICH instead of the usual AS55Q because the MICH didn't stay locked with AS55Q for some reason.

The PRCL was held by REFL33I as usual.

Also I disabled the power normalization for the error signals because it could do something bad during the Y arm is borough to the resonance.

In order to reduce the number of the glitches, PRM was slightly misaligned because I knew that the lower finesse gives fewer glitches.

ELOG V3.1.3-